OpenStudy (anonymous):

Law of Radicals Could I get a little help with radicals? (it's second semester algebra two)

OpenStudy (anonymous):

Yea ?

OpenStudy (anonymous):

\[\sqrt{6} \times \sqrt{8}\]

OpenStudy (anonymous):

multiply the insides of the radical together then take the square root

OpenStudy (anonymous):

So its \[\sqrt{48}\]

OpenStudy (anonymous):

ok well that's what i put but it came up wrong

OpenStudy (anonymous):

Because its not a perfect square, you have to do the factor tree thing ,

OpenStudy (anonymous):

you know what i mean ?

OpenStudy (anonymous):



OpenStudy (anonymous):

Ohhh so I just forgot to simplify right? so it would be \[\sqrt{3 \times 3 \times 2 \times 2 \times 2}\]

OpenStudy (anonymous):

no ,

OpenStudy (anonymous):

its \[4\sqrt{3}\]

OpenStudy (anonymous):

Oh ok I get it thanks

OpenStudy (anonymous):

your welcome.